OpenStudy (anonymous):

The perimeter of a rectangle is 50 cm and the length is 4 cm longer than the width. Write and solve the system of equations using the substitution method. A. width = 10 cm, length= 14 cm B. width = 10 1/2 cm, length= 14 1/2 cm C. width = 11 cm, length= 15 cm D. width = 23 cm, length= 27 cm

OpenStudy (anonymous):

Do you know the equation for the perimeter of a rectangle?

OpenStudy (anonymous):



OpenStudy (anonymous):

There are four sides, so to find the perimeter we just add them all up\[perimeter = length + length + width + width\]

OpenStudy (anonymous):

we can simplify this expression by using multiplication because \[width + width = 2 \times width\]

OpenStudy (anonymous):

and \[length + length = 2 \times length\]

OpenStudy (anonymous):

so the perimeter equation looks like this\[perimeter = 2 \times length + 2 \times width\]

OpenStudy (anonymous):

this problem gives us two pieces of information that we can insert into out equation for perimeter

OpenStudy (anonymous):

we are told that the perimeter = 50 and that the length is 4 more than the width so we can write the length like this \[length = width + 4\]

OpenStudy (anonymous):

we can substitute both of these pieces of information into the equation for perimeter\[perimeter = 2 \times length + 2 \times width\]

OpenStudy (anonymous):

to get:\[50 = 2 \times (width + 4) + 2 \times width\]

OpenStudy (anonymous):

By substituting length with width + 4 we reduced the equation to a single variabale which we can solve.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you handle it from here?

OpenStudy (anonymous):

yeah I had after the first two messages, but thanks for being so detailed I appreciate it

OpenStudy (anonymous):

Now how do I give you a medal?